OpenStudy (anonymous):

Suppose you have a part time job delivering packages. Your employer pays you a flat rate of 9.5 per hour. You discover that a competitor pays employees $2 per hour plus $3 per delivery. How many deliveries would the competitors employees have to have to make in four hours to earn the same pay you earn in a four hour shift? - How can you write a system of equations to model this situation? - Which method should you sue to solve the system? - How can you interpret the solution in the context of the problem

OpenStudy (anonymous):

9.5h = 2h + 3d --- sub in 4 for h because we are dealing with 4 hrs 9.5(4) = 2(4) + 3d 38 = 8 + 3d -- subtract 8 from both sides 30 = 3d -- divide by 3 on both sides 10 = d check.. 38 = 8 + 3(10) 38 = 8 + 30 38 = 38 (correct) They would have to make 10 deliveries

OpenStudy (anonymous):

thankyou very much